OpenStudy (anonymous):

Solve the inequality. Graph the solution. 2l x+1/3 l <9

OpenStudy (anonymous):

start with \[|x+\frac{1}{3}|<\frac{9}{2}\]

OpenStudy (anonymous):

then make a sandwich \[-\frac{9}{2}<x+\frac{1}{3}<\frac{9}{2}\] solve by subtracting \(\frac{1}{3}\) all the way across

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

ok im lost.

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

is this the original question \[2|x+\frac{1}{3}|<9\]?

OpenStudy (anonymous):

yes. A. -4 5/6 < x< 4 1/6 B. -4 1/3 < x < 4 1/3 C. -4 1/3 or x > 4 1/3 D. -4 5/6 or x > 4 1/6

OpenStudy (anonymous):

ok so one thing we can do is ignore the last two answers, because if you have an absolute value LESS THAN something , you are going to get an inequality that looks like this \[a<x<b\]

OpenStudy (anonymous):

when u said frac(1)(3) u mean 1/3 right?

OpenStudy (anonymous):

B was wrong

OpenStudy (anonymous):

oooh if you cannot read what i wrote refresh your browser

OpenStudy (anonymous):

so it is A?

OpenStudy (anonymous):

it is not reading the latex for some reason try refreshing and you should see it

OpenStudy (anonymous):

well, if it is not B it must be A because it is not C or D for sure

OpenStudy (anonymous):

can you read what i wrote above?

OpenStudy (anonymous):

i'll write it without the latex -9/2 < x+1/3 < 9/2 -9/2-1/3<x<9/2-1/3

OpenStudy (anonymous):

omg yes. I should've refresh it. but thank you so much!

OpenStudy (anonymous):

yw it is A, hope the method is clear

OpenStudy (anonymous):

why ru so good at math? r u a teacher?

OpenStudy (anonymous):

btw it is good to know when you start what the answer will look like if it is \[|something |<p\] it will be \[a<x<b\] but if it is \(|something |>p\) it will be \(x<a\) OR \(x>b\)

OpenStudy (anonymous):

lol, i am a car

OpenStudy (anonymous):

Okay :)